Find the modulus and the principal argument of each of the given complex numbers.
(a) 3−4i, (b) −2+i, (c) 1√ , (d) 7−i , 1+i 3 −4−3i
(e) 5(cos π/3 + i sin π/3), (f) cos 2π/3 − sin 2π
1
Expert's answer
2020-05-25T18:48:07-0400

a) "z=3-4i"

"|z|=\sqrt{3^2+4^2}=5"

"cos\theta=3/5, sin\theta=-4/5"

"\theta=-53\degree"


b) "z=-2+i"

"|z|=\sqrt{2^2+1}=\sqrt{5}"

"cos\theta=-2/\sqrt{5}, sin\theta=1/\sqrt{5}"

"\theta=153\degree"


d) "z=7-i"

"|z|=\sqrt{7^2+1}=5\sqrt{2}"

"cos\theta=\frac {7}{5\sqrt{2}}, sin\theta=-\frac {1}{5\sqrt{2}}"

"\theta=-1\degree"


"z=1+i"

"|z|=\sqrt{1+1}=\sqrt{2}"

"cos\theta=sin\theta=1/\sqrt{2}"

"\theta=45\degree"


"z=-4-3i"

"|z|=\sqrt{4^2+3^2}=5"

"cos\theta=-4/5, sin\theta=-3/5"

"\theta=-143\degree"


e) "z=5(cos \u03c0\/3 + i sin \u03c0\/3)"

"|z|=5\sqrt{cos^2\pi/3+sin^2\pi/3}=5"

"\theta=\pi/3"


f) "z=cos 2\u03c0\/3 \u2212 isin 2\u03c0\/3"

"|z|=\sqrt{cos^22\pi/3+sin^22\pi/3}=1"

"\theta=-2\pi/3"